# Scalable Language Models with Posterior Inference of Latent Thought Vectors

### v1

@@ 3 Empirical Study
Figure 4: Analysis of model scaling behavior across model size, inference steps, and latent size. Validation perplexity is shown for different configurations. Autoregressive and diffusion baselines are plotted as dashed lines.
@@ 3.1 Experimental Setup
Figure 5: Scaling Behaviors over Training Tokens and Compute. We plot the performance of LTM training runs ($N_{{\mathbf{z}}}=24$) across inference steps (16-64) and model sizes (38M-76M). Models with more inference steps demonstrate improved sample efficiency and become compute-efficient beyond certain training compute thresholds.
@@
Architectures and Training. All LTMs share similar architectures, with small, medium, and large variants using 3, 6, and 12 layers respectively. Our training was conducted on 8 H100 GPUs with an epoch batch size of 512. We employed two learning rate schedulers for dual-rate learning: fast learning schedules linearly increasing from 0.3 to 0.34, and slow learning schedules beginning at $4\times e^{-4}$ with cosine decay. Other training details are provided in Section A.2.
@@ 3.2 Scaling Behaviors
Scaling model size, inference steps, and latent size. LTMs extend traditional autoregressive models with two additional scaling dimensions: inference steps and latent size. Figure 4 shows validation perplexity across our configuration sweep.
  • Latent size: The relationship between latent size and model size represents a fundamental consideration in latent-variable models. Larger latent sizes lead to reduced sample efficiency (measured by validation perplexity) and decreased training stability in our dual-rate learning framework. However, since latent size minimally impacts trFLOPs/tok, we can optimize it independently from inference steps.
  • Inference steps vs model size: Performance improvements from inference steps become apparent starting from 16 steps. At fixed latent and model size, performance improves with additional inference steps until reaching a plateau at approximately 64 steps. Conversely, at fixed latent size and inference steps, model size has minimal impact, likely because attention layers’ contribution has not yet overtaken that of embedding layers at this scale.
Considering both trFLOPs/tok and training dynamics, these scaling dimensions form a hierarchical structure for model design, as illustrated in Figure 1.
Inference steps drive sample and compute efficiency. When extrapolating scaling properties to larger training compute regimes, converged performance becomes less relevant for model selection. As demonstrated by Kaplan et al. (2020), training larger models without reaching convergence proves more compute-efficient than training smaller models to convergence. Figure 5 shows that LTMs possess similar properties: models with more inference steps achieve greater sample efficiency and become more compute-efficient beyond certain thresholds of training compute. The minimal influence of model size on these curves likely stems from embedding layers’ computation remaining comparable to attention layers at this scale.
@@ 3.3 Comparison with Existing Language Models
LTM-Small (steps$=16$) $38$M $4.07$G $7$B $\leq 34.71$ $\leq 18.87$ $\leq 23.59$ $\leq 19.31$ $\leq 34.76$ $\leq 22.73$ $\leq 21.67$
LTM-Medium (steps$=16$) $51$M $5.52$G $5.2$B $\leq 32.06$ $\leq 17.39$ $\leq 25.16$ $\leq 17.32$ $\leq 27.89$ $\leq 20.45$ $\leq 19.22$
LTM-Large (steps$=64$) $76$M $32.2$G $0.9$B $\leq\mathbf{12.02}$ $\leq\mathbf{8.02}$ $\leq\mathbf{9.63}$ $\leq\mathbf{8.47}$ $\leq\mathbf{10.55}$ $\leq\mathbf{9.39}$ $\leq\mathbf{8.91}$
Pretraining Perplexity. LTMs’ perplexities on OWT validation set are marked in Figure 4. The inference compute for this evaluation is close to trFLOPs/tok, expect that there is no slow learning. Trained with equivalent trFLOPs/tok as GPT-2-L, LTM-Medium performs slightly better, with only 10% parameters. The model size can be further reduced to 38M in LTM-Small, maintaining similar performance. LTM-Large achieves state-of-the-art validation perplexity: 5.58 even if it is only trained with $3$B tokens.
@@
We then explore LTMs few-shot in-context learning capability, which traditionally emerges only at GPT-3 scale (Brown et al., 2020). Using randomly sampled training examples as in-context demonstrations, we find that LTMs exhibit this capability even in our most lightweight configuration (38M parameters). As shown in Figure 6, LTM-Small with 5-shot demonstrations surpasses the baselines from Li et al. (2022) that incorporates finetuning or test-time search. Increased model size further improves both zero-shot and few-shot performance. Motivated by the hypothesis that a more expressive latent space enables stronger abstract reasoning, we tested an LTM-Large variant with $192$ latent thought vectors, which achieves the best performance. Additional experiment details are included in Appendix A.3.
LTMs’ few-shot learning capability differs fundamentally from related approaches. Unlike autoregressive models (Brown et al., 2020), LTMs use gradient-based inference for latent thought vectors, enabling few-shot learning at much smaller scales. This suggests more efficient pattern discovery at abstract levels. The emergent nature of this capability contrasts with meta-learning methods that require bi-level optimization on downstream tasks (Finn et al., 2017; Yoon et al., 2018) — LTMs achieve few-shot learning directly within the context window without specialized training.
@@
While GPT-2 requires nucleus sampling to achieve comparable performance with diffusion models, LTMs outperform both approaches using standard multinomial sampling. As shown in LABEL:table:conditional_gen, LTMs maintain nearly equivalent performance even with greedy decoding, suggesting that the per-token distribution conditioned on latent thought vectors, $p_{\beta}(x^{(n)}|{\mathbf{z}},{\mathbf{x}}^{(<n)})$, is highly concentrated. We include additional samples in Appendix A.5.
@@
As both terms are intractable, alternative metrics have been proposed: Dieleman et al. (2022) introduce generative perplexity (Gen PPL), which approximates $p_{\rm{data}}$ in the first term using a larger language model, while Zheng et al. (2024) propose token-level entropy to approximate the second term and detect mode collapse. We use LTM-XLarge as the proxy for $p_{\rm{data}}$ to calculate the Gen PPL.
LABEL:table:uncond_gen presents the results. While SEDD-M achieves a Gen PPL of 32.63 with 1024 sampling steps and an entropy of 5.27, we follow Zheng et al. (2024)’s recommendation to consider only baselines with entropy exceeding 5.6. Under these criteria, LTM-Large achieves performance comparable to GPT-2-Large on both metrics while providing a 5×5$\times$5 × faster sampling speed. Experiment details can be found in Appendix A.3, with additional samples in Appendix A.4.
@@ 3.4 Ablation Studies
We explore inference strategies for LTMs. Our VAE baseline, which employs an identical decoder and a 12-layer encoder with full attention, suffers from posterior collapse, resulting in repetitive prior samples and low entropy distributions. While implementing Langevin sampling with LTMs using the same decoder helps mitigate posterior collapse, it produces lower quality generations compared to variational learning approaches.
@@ 4 Related work
Variational Bayes Language Modeling. In the context of language models, Bowman et al. (2016) introduced a variational autoencoder for text generation. Building on this, Xu & Durrett (2018) proposed the use of von Mises-Fisher distribution in VAEs. Li et al. (2020) presented OPTIMUS, a large-scale pretrained deep latent variable model for natural language. Pang & Wu (2021); Yu et al. (2022) studied VAE for language modeling with learnable prior model.
Large Language Models with Explicit Latent Space. Zelikman et al. (2022); Hu et al. (2023); Hoffman et al. (2024) repurpose token-level LLMs to generate latent chains of thought. Hao et al. (2024) repurpose the hidden state of Transformers as continuous latent space. They are all post-training methods that demonstrate the advantages of explicit latent learning. Concurrent to our work, The et al. (2024) train generative models for the latent embedding of a pretrained auto-encoder.
@@
5 Discussion on prior model in latent space
Our current work assumes a simple Gaussian prior model for the latent thought vectors. The only structural design we employ is to assume separate sets of thought vectors that cross-attend to different layers of Transformer decoder. While such a simple prior model is a suitable starting point for initial systematic investigation, especially the scaling properties, much can be gained by imposing more structured and learnable prior models. For instance, logical deduction and language of thoughts (Fodor, 1975) may be modeled by a latent reasoning decoder model that generates a chain of latent thought vectors in the lifted latent space, which is then rendered into ground tokens by a separate grounding decoder. Although we may impose general structures on the prior model, we want to emphasize that no hand-designed logical or grammar rules are imposed. Instead, the hope is that they may emerge naturally from learning such structured top-down probabilistic models.
The importance of a learnable prior model in a lifted latent space cannot be overstated. Such explicit structured priors offer several crucial advantages: First, they establish an internal fixed-form representation of knowledge, transforming posterior inference into a process of parsing, formalization, and understanding. Second, they enable top-down modeling of abstraction hierarchies, effectively separating concepts from their instantiations and variables from their assigned values. Third, they support top-down modeling of compositional hierarchies, capturing part-whole relationships at various scales. Finally, they hold significant potential for facilitating reasoning, search, planning, and reinforcement learning within a compact abstract space. These capabilities stand in stark contrast to prevalent AI models that operate directly on ground tokens or signals (such as pixel intensities in images and videos). While posterior inference in lifted latent spaces can be computationally intensive, the benefits — including improved data efficiency, enhanced extrapolation and invention capabilities, and greater interpretability and controllability — shall justify this computational heavy lifting.
@@ 6 Conclusion
In this paper, we introduce Latent-Thought Language Models (LTMs), which incorporate explicit latent thought vectors that follow explicit prior models in latent space. We develop a novel dual-rate optimization algorithm for training these models and conduct extensive empirical investigations of their properties, with particular focus on scaling behaviors along inference steps and latent dimensionality. Our approach draws inspiration from cognitive science theories, including declarative-procedural memory systems, the language of thought hypothesis, and complementary learning systems. Our work lays the groundwork for further development of more structured prior models in the lifted latent space for the purpose of reasoning and planning.
@@ Acknowledgment
We thank Ruiqi Gao and Kevin Murphy for insightful discussions and valuable suggestions. Y. W. was partially supported by NSF DMS-2015577, NSF DMS-2415226, and a gift fund from Amazon. We gratefully acknowledge the support of Lambda, Inc. for providing compute resources for this project.
@@ A.2 Training Details
We train all models using a “slow” learning rate of $4\times e^{-4}$ followed by cosine decay schedule to $4\times e^{-5}$. We also apply a linear warmup schedule to the first 1000 iterations, and clip the gradient norm to 1 during training. For the “fast” learning rate, we start from $0.3$ and linearly increases to $0.34$.
For both global parameters and latent thought vectors update, we use AdamW optimizer (Loshchilov, 2017) with $\beta_{1}=0.9$, and $\beta_{2}=0.95$.

### v2

@@ 3.1 Experimental Setup
Figure 4: Scaling behaviors over training tokens and compute. We plot the performance of LTM training runs across inference steps ($T_{\text{fast}}=$16-64), latent size ($N_{{\mathbf{z}}}=$24-96) and model sizes (38M-76M). Models with more inference steps demonstrate improved sample efficiency and become compute-efficient beyond certain training compute thresholds.
@@
Architectures and Training. All LTMs share similar architectures, with small, medium, and large variants using 3, 6, and 12 layers respectively. Our training was conducted on 8 H100 GPUs with an epoch batch size of 512. We employed two learning rate schedulers for dual-rate learning: fast learning schedules linearly increasing from 0.3 to 0.34, and slow learning schedules beginning at $4\times 10^{-4}$ with cosine decay. Other training details are provided in Section A.2.
@@ 3.2 Scaling Behaviors
Scaling model size, inference steps, and latent size. LTMs extend traditional autoregressive models with two additional design axes: inference steps and latent size. Fig. 1 shows validation perplexity across our configuration sweep.
  • Latent size: More latent thought vectors improve performance across all model sizes and inference step configurations. The 76M parameter models show clear performance gains when increasing from $N_{\mathbf{z}}=24$ to to $N_{\mathbf{z}}=96$, indicating that latent dimensionality serves as an effective scaling dimension for LTMs.
  • Inference steps vs model size: Performance improvements from inference steps become apparent starting from 16 steps to 128 steps. For larger steps, we find that scheduling the fast learning rate helps for stable training, In particular, we adopt a cosine decay scheduler. Conversely, at fixed latent size and inference steps, model size has minimal impact, likely because attention layers’ contribution has not yet overtaken that of embedding layers at this scale.
Inference steps drive sample and compute efficiency. When extrapolating scaling properties to larger training compute regimes, converged performance becomes less relevant for model selection. As demonstrated by Kaplan et al. (2020), training larger models without reaching convergence proves more compute-efficient than training smaller models to convergence. Fig. 4 shows that LTMs possess similar properties: models with more inference steps achieve greater sample efficiency and become more compute-efficient beyond certain thresholds of training compute. Additionally, larger latent sizes ($N_{\mathbf{z}}=48,96$) further enhance both sample and compute efficiency when combined with more inference steps. The minimal influence of model size on these curves likely stems from embedding layers’ computation remaining comparable to attention layers at this scale.
@@ 3.3 Comparison with Existing Language Models
LTM-Small ($T_{\text{fast}}=16$) $38$M $4.07$G $7$B $\leq 34.71$ $\leq 18.87$ $\leq 23.59$ $\leq 19.31$ $\leq 34.76$ $\leq 22.73$ $\leq 21.67$
LTM-Medium ($T_{\text{fast}}=16$) $51$M $5.52$G $5.2$B $\leq 32.06$ $\leq 17.39$ $\leq 25.16$ $\leq 17.32$ $\leq 27.89$ $\leq 20.45$ $\leq 19.22$
LTM-Large ($T_{\text{fast}}=64$) $76$M $32.2$G $0.9$B $\leq\mathbf{4.43}$ $\leq\mathbf{3.66}$ $\leq\mathbf{3.92}$ $\leq\mathbf{3.48}$ $\leq\mathbf{4.56}$ $\leq\mathbf{3.87}$ $\leq\mathbf{3.54}$
Pretraining Perplexity. LTMs’ perplexities on OWT validation set are marked in Fig. 1. The inference-time compute for this evaluation is close to trFLOPs/tok, except that there is no slow learning. Trained with equivalent trFLOPs/tok as GPT-2-Large, LTM-Medium performs slightly better, with only 10% parameters. The model size can be further reduced to 38M, as in LTM-Small, without compromising much performance. LTM-Large achieves state-of-the-art validation perplexity: 3.05 even if it is only trained with $3$B tokens. While more inference steps could yield higher sample efficiency, and better perplexity we choose LTM-Large as it provides a favorable tradeoff between inference speed and sample efficiency.
@@
We then explore LTMs few-shot in-context learning capability, which traditionally emerges only at GPT-3 scale (Brown et al., 2020). Using randomly sampled training examples as in-context demonstrations, we find that LTMs exhibit this capability even in our most lightweight configuration (38M parameters). As shown in Fig. 5, LTM-Small with 5-shot demonstrations surpasses the baselines from Li et al. (2022) that incorporates finetuning or test-time search. Increased model size further improves both zero-shot and few-shot performance. Motivated by the hypothesis that a more expressive latent space enables stronger abstract reasoning, we tested an LTM-Large variant with $192$ latent thought vectors, which achieves the best performance. Additional experiment details are included in Section A.3.
LTMs’ few-shot learning capability differs fundamentally from related approaches. Unlike autoregressive models (Brown et al., 2020), LTMs use gradient-based inference for latent thought vectors, enabling few-shot learning at much smaller model scales. This suggests more efficient pattern discovery at abstract levels. The emergent nature of this capability contrasts with meta-learning via bi-level optimization on downstream tasks (Finn et al., 2017; Yoon et al., 2018) — LTMs achieve few-shot learning directly within the context window without specialized training.
@@
While GPT-2 requires nucleus sampling to achieve comparable performance with diffusion models, LTMs outperform both approaches using standard multinomial sampling. As shown in LABEL:table:conditional_gen, LTMs maintain nearly equivalent performance even with greedy decoding, suggesting that the per-token distribution conditioned on latent thought vectors, $p_{\beta}(x^{(n)}|{\mathbf{z}},{\mathbf{x}}^{(<n)})$, is highly concentrated. We include additional samples in  Section A.5.
@@
As both terms are intractable, alternative metrics have been proposed: Dieleman et al. (2022) introduce generative perplexity (Gen PPL), which approximates $p_{\rm{data}}$ in the first term using a larger language model, while Zheng et al. (2024) propose token-level entropy to approximate the second term and detect mode collapse. We use GPT-2-XL as the proxy for $p_{\rm{data}}$ to calculate the Gen PPL.
LABEL:table:uncond_gen presents the results. While SEDD-M achieves a Gen PPL of 32.63 with 1024 sampling steps and an entropy of 5.27, we follow Zheng et al. (2024)’s recommendation to consider only baselines with entropy exceeding 5.6. Under these criteria, LTM-Large achieves performance comparable to GPT-2-Large on both metrics while providing a 5×5$\times$5 × faster sampling speed. Experiment details can be found in Section A.3, with additional samples in Section A.4.
@@ 3.4 Ablation Studies
We explore inference strategies for LTMs. Our VAE baseline, which employs an identical decoder and a 12-layer encoder with full attention, suffers from posterior collapse, resulting in repetitive prior samples and low entropy distributions. While implementing Langevin sampling with LTMs using the same decoder helps mitigate posterior collapse, it produces lower quality generations compared to the variational Bayes learning approach.
@@
3.5 Probing Results on Latent Thought Vectors
We investigate how semantic information distributes hierarchically across LTMs’ layers through progressive reconstruction experiments, where we evaluate reconstruction accuracy by progressively including layers of latent thought vectors from bottom to top.
The study in Fig. 10 reveals that LTMs process information in a layered fashion, with different model sizes showing distinct hierarchical patterns. For the 12-layer LTM model with 96 latent thought vectors, we observe distributed information processing with steady increases in reconstruction accuracy through bottom and middle layers (1-8), reaching approximately 65% accuracy. This is followed by crucial synthesis at top layers (9-10), where accuracy jumps dramatically to over 95%. The case study in Fig. 11 demonstrates this clear semantic progression. Bottom layers produce scattered, disconnected terms, middle layers develop structural coherence with emerging phrases and descriptive elements, while top layers achieve complete semantic integration and perfect reconstruction. This hierarchical organization reveals distinctive “synthesis layers” in the top of the network that integrate information from earlier layers, showing how LTMs encode and process semantic information through the layered thought vectors. See Appendix B for more details.
4 Limitations: Prior and Reward
Learnable Structured Prior Models. Our current work assumes a simple Gaussian prior model for the latent thought vectors. The only structural design we employ is to assume separate sets of thought vectors that cross-attend to different layers of Transformer decoder. While such a simple prior model is a suitable starting point for initial systematic investigation, much can be gained by imposing a more structured and learnable prior model with more interpretable latents, $p_{\alpha}({\mathbf{z}})$. For instance, language of thoughts (Fodor, 1975) may be modeled by a latent reasoning model that generates a chain of latent thought vectors in the latent space, transforming posterior inference into a process of parsing, formalization, compression, and understanding.
Reward or Verifier Models in Latent Space. Our model currently lacks a reward model or verifier model defined in the latent space, $p_{\gamma}(r|{\mathbf{z}})$, which can be used to guide the optimization of ${\mathbf{z}}$ as a form of inference-time computation for reasoning. In our recent work on latent plan transformer models, we have applied such models to offline reinforcement learning (Kong et al., 2024b) and online optimization for molecule design (Kong et al., 2024a).
@@ 5 Related Work
Variational Bayes Language Modeling. Bowman et al. (2016) introduce a variational autoencoder for text generation. Building on this, Xu & Durrett (2018) propose the use of von Mises-Fisher distribution in VAEs. Li et al. (2020) present OPTIMUS, a large-scale pretrained deep latent variable model for natural language. Pang & Wu (2021); Yu et al. (2022); Xu et al. (2023) study language modeling with learnable prior model.
Large Language Models with Explicit Latent Space. Zelikman et al. (2022); Hu et al. (2023); Phan et al. (2023) repurpose token-level LLMs to generate latent chains of thought. Hao et al. (2024) repurpose the hidden state of Transformers as continuous latent space. They are all post-training methods that demonstrate the advantages of explicit latent learning. Concurrent to our work, The et al. (2024) train generative models for the latent embedding of a pretrained auto-encoder.
@@ 6 Conclusion
In this paper, we introduce Latent Thought Models (LTMs), which incorporate explicit latent thought vectors that follow explicit prior models in latent space. We develop a novel dual-rate optimization algorithm for training these models and conduct extensive empirical investigations on their properties, with particular focus on scaling behaviors along inference steps and latent dimensionality. Our approach draws inspiration from cognitive science theories, including declarative-procedural memory systems, the language of thought hypothesis, and complementary learning systems. Our work lays the groundwork for further development of more structured and interpretable prior models and reward-verifier models in the latent space for the purpose of reasoning and planning.
@@ Acknowledgment
We thank Ruiqi Gao and Kevin Murphy for insightful discussions and valuable suggestions. Y. W. was partially supported by NSF DMS-2015577, NSF DMS-2415226, and a gift fund from Amazon. We gratefully acknowledge the support of Lambda, Inc. for providing the compute for this project.
Impact Statement
Our paper investigates a new model class for language modeling with explicit latent thought vectors and inference-time computation. This model class has the potential to learn more explicit internal representations and enable more explicit reasoning and planning based on such representations.
@@ A.2 Training Details
We train all models using a “slow” learning rate of $4\times 10^{-4}$ followed by cosine decay schedule to $4\times 10^{-5}$. We also apply a linear warmup schedule to the first 1000 iterations, and clip the gradient norm to 1 during training. For the “fast” learning rate, we start from $0.3$ and linearly increases to $0.34$.
We use AdamW optimizer (Loshchilov, 2017) with $\beta_{1}=0.9$, and $\beta_{2}=0.95$ to update the global parameters. We use Adam to update the latent thought vectors without introducing additional inductive bias in the optimization.
@@ A.5 Samples for Conditional Generation
Appendix B Probing the Latent Thought Vectors
To understand how LTMs hierarchically encode information, we evaluate reconstruction accuracy by progressively including layers of latent thought vectors from bottom to top across 200 samples from the OpenWebText validation set. We test two model configurations shown in Fig. 10: LTM-Medium (6-layer, 24 latent vectors with 4 per layer) and LTM-Large (12-layer, 96 latent vectors with 8 per layer), measuring how reconstruction accuracy improves as we incrementally include more layers during text generation. Additionally, we present a detailed case study in Fig. 11 that demonstrates the specific reconstruction patterns emerging at each layer of latent thought vectors.
B.1 Progressive Layer Inclusion
Figure 10: Left: 6-layer LTM-Medium with with 24 latent vectors (4 per layer). Right: 12-layer LTM-Large with 96 latent vectors (8 per layer). Distribution of Reconstruction Accuracy with Progressive Layer Inclusion for LTM models. The plots show how reconstruction accuracy improves as layers are progressively included from bottom to top, measured across 200 sequences from OpenWebText validation set. (a) 6-layer LTM-Medium shows gradual improvement through layers 1-5 ($\sim$55accuracy) followed by a sharp jump at layer 6 to complete reconstruction. (b) 12-layer LTM-Large demonstrates more distributed information processing with steady increases through layers 1-8 ($\sim$65%), followed by crucial synthesis at layers 9-10, reaching $>$95% accuracy. This reveals the hierarchical nature of LTMs’ latent representations, with deeper models distributing information more gradually across layers and featuring distinctive “synthesis layers” that integrate information from earlier representations.
B.2 Case Study
Progressive Inclusion of Latent Thought Vectors ($z$) Using Layers 1-3 Only (22% Accuracy): The … … to its … … … … … … … … … … of … … . The … … to its surface, … … … … … … … sea … … … … … … … … … … … … … … … … … … … … … … … … … … … … . … and I sea … … … … … … … … … … in … . The … … of … … … … … … … … … with … . … … … … … … … … … in the … … … … and … … … … … to … . the … … … … … its … the … … surface. … … … … … … … … as … a … of the sea … of … … a sea of … … … … … . … … … … to be … … … … of … … with the … … and … … the … … … … and the … … … of Using Layers 1-6 Only (30% Accuracy): The … … through way … the … … … along each … … … … … to … … while … … … beneath sea and … into … sea of … … … carried … … … … … … of … … … Wild … … … and in … of … … … by … … blue … … … … … … secluded … … … … … … … tide. A … waters had … a … … of … … … … … … with … . … … … … … … … … … … between … … … … . … … the … began to … … … … … … of light across … … surface of the ocean. … the … … … boat … as … more than a … … the … horizon, … a procession of … seabirds in its wake. The … … to exist in a perpetual … of … … by … tide, and season… … … … to … … and the … of … . Using Layers 1-9 Only (65% Accuracy): … … its way along the … cliffs, revealing new … with … turn … … mist … to … … … … … between sea and … into a gradient of silvery … . … air carried … distant … of gulls and … … … of … against stone. … heather … … … in … of purple, … interrupted by … … yellow … gorse flowers. … where the trail … toward … … cove, its crescent … … … visible only at low tide. The … waters had … a … of tide … , each a … … … with life. … seawater … , … water, while tiny … scuttled between … … … … . Overhead, the sun began to … through the … , casting … of light across … undulating surface of the ocean. … the … , a … … appeared as little more than a … against the … horizon … trailing a procession of opportunistic seabirds in its wake. The … … … exist … … … … … … tide, and season… somehow unchanged, indifferent to … concerns and the … … … . Using Layers 1-10 Only (99% Accuracy): … coastal path wound its way along the rugged cliffs, revealing new vistas with each turn. Morning mist clung to the landscape, softening the boundary between sea and sky into a gradient of silvery blues. Salt-laden air carried the distant cries of gulls and the rhythmic percussion of waves against stone. Wild heather painted the hillsides in swathes of purple, occasionally interrupted by the defiant yellow of gorse flowers. I paused where the trail dipped toward a secluded cove, its crescent of golden sand visible only at low tide. The receding waters had revealed a tapestry of tide pools, each a miniature universe teeming with life. Emerald seaweed swayed in crystalline water, while tiny crabs scuttled between barnacle-encrusted rocks. Overhead, the sun began to burn through the haze, casting diamonds of light across the undulating surface of the ocean. In the distance, a fishing boat appeared as little more than a silhouette against the brightening horizon, trailing a procession of opportunistic seabirds in its wake. The landscape seemed to exist in a perpetual state of change—shaped by wind, tide, and season—yet somehow timeless, indifferent to human concerns and the passage of years. Using All Layers (1-12) (100% Accuracy): The coastal path wound its way along the rugged cliffs, revealing new vistas with each turn. Morning mist clung to the landscape, softening the boundary between sea and sky into a gradient of silvery blues. Salt-laden air carried the distant cries of gulls and the rhythmic percussion of waves against stone…
Figure 11: Progressive reconstruction of text using latent thought vectors from a 12-layer LTM. This figure displays only the correctly reconstructed words at each layer, showing how text accuracy improves as more layers are included. Dots (…) represent incorrect or missing words. Color coding: purple for partial reconstructions and orange for near-complete or complete reconstructions. At layer 0-3 (22% accuracy), only scattered words match the original. By layer 0-6 (30%), more structural elements emerge, including some phrases about the ocean and landscape. Layer 0-9 (65%) shows substantial improvement with coherent phrases and key descriptive elements. Complete accuracy (100%) is achieved with all 12 layers. This progression demonstrates how semantic information is hierarchically distributed across the model’s latent space.